A ship traveled at an average rate of 25 miles per hour going west. It then traveled at an average rate of 19 miles per hour heading north. If the ship traveled a total of 145 miles in 7 hours, how many miles were traveled heading west?
Answer by ankor@dixie-net.com(22740)

A ship traveled at an average rate of 25 miles per hour going west.
It then traveled at an average rate of 19 miles per hour heading north.
If the ship traveled a total of 145 miles in 7 hours, how many miles were traveled heading west?
:
Find how long the ship traveled at each speed
let w = time heading west at 25 mph
the total time is 7 hrs, therefore
(7-w) = time heading north at 19 mph
:
A distance equation; dist = speed * time
25w + 19(7-w) = 145
25w + 133 - 19w = 145
25w - 19w = 145 - 133
6w = 12
w = 12/6
w = 2 hrs traveling west
therefore
2 * 25 = mile going west
:
:
Confirm this solution, 7 - 2 = 5 hrs going north
19 * 5 = 95mi
25 * 2 = 50mi
--------------
total dist 145 mi